OpenStudy (anonymous):

(4y^a - 6x^b)(4y^a + 6x^b)

OpenStudy (mrnood):

this is in the form (a+b)(a-b) can you see that? (a+b)(a-b) = a^2 - b^2 so substitute the terms for a and b then do the 'squaring' then simplify

OpenStudy (anonymous):

so 2a-2b ??

OpenStudy (mrnood):

^2 means 'squared'

OpenStudy (mrnood):

I may have confused oyu by using a & b my equation is a 'general form. look at YOUR equation and work out wht the equivalent of a & b are in my equaiton

OpenStudy (mrnood):

(m+n)(m-n) = m^2 - n^2 That avoids the confusion with the a & b in your equation So look at the FORM of your equation, and find an equivalent to m & n